OpenStudy (chaotic_butterflies):

2^(3x+3) = 1 Solve for x I've simplified the problem from 8^(x-1) = 1/64 down to the problem which you see above, and I don't know what further step to take. I've been working with logs in this chapter

OpenStudy (aravindg):

Take log on both sides

OpenStudy (aravindg):

Log a^x =x loga

OpenStudy (inkyvoyd):

there are two ways to do this. the first way will help you understand logs. the second way is what @AravindG said. Think this way: 2 to the power of what equals 1?

OpenStudy (chaotic_butterflies):

I can't just get that, and what you both are hinting at sounds familiar, but I'm still unsure what to do. If I take the logs of both sides, would it look like this: 2 log(3x+3) = log (1) ?

OpenStudy (inkyvoyd):

your 2 and 3x+3 are switched

OpenStudy (chaotic_butterflies):

Ah, but is the right side of my equation correct? and do I just simplify from there?

OpenStudy (inkyvoyd):

yes.

OpenStudy (chaotic_butterflies):

so after I simplify log(2), do I multiply it by the three or do I add it?

OpenStudy (inkyvoyd):

might want to show your work; not sure i can follow

OpenStudy (aravindg):

Log a^x =x loga

OpenStudy (chaotic_butterflies):

3x + 3 log (2) = log(1) 3x + 3 * 0.301 = 0

OpenStudy (chaotic_butterflies):

is that how I should go about it?

OpenStudy (inkyvoyd):

well, you're missing some parenthesis. Let's not simplify the log of 2, rather, recognize that the logarithm of 1 in any base is always 0... (why is this? isn't a^0 always equal to 1)?...

OpenStudy (chaotic_butterflies):

I did that on the right side, I don't know what to do for the left side. How do I address the log of two in relation to the rest of the right side

OpenStudy (inkyvoyd):

(3x + 3)log (2) = log(1) (3x + 3)log (2) = 0 log of 2 is just a number. let's divide both sides by log(2) and then by 3... what do we get?

OpenStudy (chaotic_butterflies):

wouldn't I just be left with 3x+3 = 0 then?

OpenStudy (inkyvoyd):

well, if you didn't divide by 3... either way you can solve for x I am sure now ;)

OpenStudy (chaotic_butterflies):

Ah, thanks!

OpenStudy (inkyvoyd):

No problem - advice for these problems, always check your answer(s) to make sure they satisfy the original equation, because you can get extraneous solutions... in this case you won't, but just be careful

OpenStudy (chaotic_butterflies):

Okie dokie